Above what Fe2+ concentration will Fe(OH)2 precipitate from a buffer solution that has a pH of 8.36 ?

The
1
Expert's answer
2020-04-28T12:25:21-0400

We know that, Ksp of Fe(OH)2="4.87\times10^{-17}"

pOH= 14-8.36=5.64

"[OH^{-1}]=-log(5.64)=2.291\times10^{-6}"


"[Fe^{2+}][OH^{-1}]=4.87\times10^{-17}"


"[Fe^{2+}][2.291\times10^{-6}]=4.87\times10^{-17}"


"[Fe^{2+}]=2.12 \times 10^{-11}"


Above this it will make buffer